OpenStudy (anonymous):

Find cot θ if csc θ = sqr5/2 and tan θ > 0. also Verify the identity. cos 4u = cos22u - sin^2 (2u) Find the exact value by using a half-angle identity. cos(pi/12)

OpenStudy (paxpolaris):

 csc positive means sin positive sin and tan are positive in 1st Quadrant

OpenStudy (paxpolaris):



OpenStudy (paxpolaris):

\[\Large \cot \theta = \frac AO=\frac12\]

OpenStudy (anonymous):

sorry i have the first one its the other two i need help with

OpenStudy (aum):

4u = 2 * 2u Use the double angle identity of cosine.

OpenStudy (anonymous):

im still confused

OpenStudy (aum):

\[ \cos(2a) = \cos^2(a) - \sin^2(a) \\ \cos(4u) = \cos(2*2u) = \cos^2(2u) - \sin^2(2u) \\ \]

OpenStudy (anonymous):

the actual problem is cos4u=cos^2(2u)-sin^(2u)

OpenStudy (aum):

The first line in the previous reply is a well-known trig identity. The second line in the previous reply answers your problem.

OpenStudy (anonymous):

oh ok i see

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

now i just need help with the second one

OpenStudy (aum):

\[ \cos(2a) = 2\cos^2(a) - 1 \\ \text{Let }a = \frac{\pi}{12} \\ \cos(2* \frac{\pi}{12}) = 2\cos^2(\frac{\pi}{12}) - 1 \\ \cos(\frac{\pi}{6}) = 2\cos^2(\frac{\pi}{12}) - 1 \\ \frac{\sqrt{3}}{2} = 2\cos^2(\frac{\pi}{12}) - 1 \\ \frac{\sqrt{3}}{2} + 1 = 2\cos^2(\frac{\pi}{12}) \\ \frac{\sqrt{3}+2}{2} = 2\cos^2(\frac{\pi}{12}) \\ \frac{\sqrt{3}+2}{4} = \cos^2(\frac{\pi}{12}) \\ \cos^2(\frac{\pi}{12}) = \frac{\sqrt{3}+2}{4} \\ \cos(\frac{\pi}{12}) = \frac{\sqrt{\sqrt{3}+2}}{2} \\ \]

OpenStudy (anonymous):

thank you dude

OpenStudy (anonymous):

do you think you have a little more time because the review only has 3 more questions?

OpenStudy (anonymous):

what is tan(x) if cos(x)=1/4 and sin(x) <0 write with only sin and cos: sin(3x)-cosx and verify the identity sin(x+pi/2)=cos(x)

OpenStudy (aum):

Use the same method shown for the previous problem. Draw a triangle. Use the definition of cosine to mark the proper sides with 1 and 4. Find the third side. Use definition of tan(x) to find its value. Since tan = sin/cos and sin < 0 and cos > 0, tan < 0.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

ok i got sqr 15